Convert the following ratios to percentages
3:4 
Convert the following ratios to percentages
3:4 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

2:3= 3/4 = 3/4 ×100/100 = 3/4 ×100% = 0.75 ×100% = 75 %​

Thanks